# KernelSHAP-IQ: Weighted Least Square Optimization for Shapley Interactions

### v1

@@ 3 Weighted Least Square Optimization for Shapley Interactions
3.1 Optimal Approximations via Shapley Interactions
Similar to the SV, we construct a $k$-additive approximation of $\nu(T)$ as the sum of $k$-SII values contained in $T$.
Definition 3.1 ($k$-additive Approximation).
For $T\subseteq N$, the $k$-additive approximation is $\hat{\nu}_{k}(T):=\sum_{S\subseteq T}^{1\leq|S|\leq k}\Phi_{k}(S)$.
Clearly, $\hat{\nu}_{k}$ includes $k$-SII and thereby SII up to order $k$. Representing $\hat{\nu}_{k}$ with SII reveals a surprising recursion.
@@ 3.1 Optimal Approximations via Shapley Interactions
For SII, it holds
$\hat{\nu}_{k}(T)=\hat{\nu}_{k-1}(T)+\sum\nolimits_{S\subseteq N}^{|S|=k}\phi^{\text{SII}}(S)\lambda(|S|,|S\cap T|),$
@@
It holds $\lambda(k,k)=0$ for $k>1$, $\hat{\nu}_{1}(T)=\sum_{i\in T}\phi^{\text{SV}}(i)$ and $\hat{\nu}_{2}(T)=\hat{\nu}_{1}(T)-\frac{1}{2}\sum_{{ij}\subseteq N}^{|ij\cap T|=1}\phi^{\text{SII}}(ij)$.
Increasing $k$ successively builds on the SV-based approximation $\hat{\nu}_{1}(T)$ by including all SII of order $k$, i.e. $\binom{n}{k}$ components, which yields a complexity-accuracy trade-off. For $k=n$, we have $\hat{\nu}_{n}(T)=\nu(T)$, which directly follows from a game-theoretic result (Grabisch et al., 2000).
@@
For any $T\subseteq N$ it holds $\nu(T)=\hat{\nu}_{n}(T)$.
In this case, $n$-SII values, which construct $\hat{\nu}_{n}$, are the Möbius transform (Bordt & von Luxburg, 2023), which is related to the well-known functional ANOVA decomposition (Hooker, 2004) from statistics, if $\nu$ are conditional expectations (Herren & Hahn, 2022).
We now link SII of order $k$, $\phi^{\text{SII}}_{k}$, to an optimal $k$-additive approximation of $\nu$ via $\hat{\nu}_{k}$. First, we introduce matrix notations, where the rows correspond to subsets $T\subseteq N$ and the columns to interactions $S$ of order $k$. For vectors $\mathbf{y}_{k}\in\mathbb{R}^{2^{n}}$, $\phi^{\text{SII}}_{k}\in\mathbb{R}^{\binom{n}{k}}$ and matrix $\mathbf{X}_{k}\in\mathbb{R}^{2^{n}\times\binom{n}{k}}$, we let $(\mathbf{y}_{k})_{T}:=\nu(T)-\hat{\nu}_{k-1}(T)$ and obtain by Proposition 3.2
$\mathbf{y}_{k}=\mathbf{X}_{k}\phi^{\text{SII}}_{k}\text{ with }(\mathbf{X}_{k})_{TS}:=\lambda(|S|,|T\cap S|).$
Introducing a diagonal weight matrix $\mathbf{W}_{k}\in\mathbb{R}^{2^{n}\times 2^{n}}$, we aim to characterize $\phi^{\text{SII}}_{k}$ as the solution to a WLS problem
@@
where $\|\cdot\|_{2}$ is the Euclidean norm. The solution is given as
@@
We thus introduce the weights $\mu_{k}$ with $\mu_{\infty}\gg 1$ as
@@
The weights $\mu_{k}$ appear naturally from $\phi^{\text{SII}}(S)=\sum_{T\subseteq N}\nu(T)\omega_{s}(t,|T\cap S|)$ (Fumagalli et al., 2023) as the common factors $\mu_{k}(t)\propto\omega_{s}(t,0),\dots\omega_{s}(t,k)$ for $k\leq t\leq n-k$. Since $(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}\in\mathbb{R}^{\binom{n}{k}\times\binom{n}{k}}$ does not depend on $T$, it holds that $\mu_{k}$ controls the final weight in $\mathbf{X}_{k}^{T}\mathbf{W}_{k}$ for all interactions $S$.
$\mu_{1}$ are the KernelSHAP weights and the SV is the solution of the WLS problem for $\mu_{\infty}\to\infty$, which justifies KernelSHAP.
@@
Note that, for $k=1$, this limit ensures the efficiency condition as $(\mathbf{X}_{1}\phi)_{N}=(\mathbf{y}_{1})_{N}=\sum_{i\in N}\phi(i)$. For pairwise SII, we now present a novel representation akin to Theorem 3.6.
@@
While $\mu_{\infty}\to\infty$ corresponds to a constraint for $k=1$ that ensures efficiency, its behavior for $k=2$ is less clear. Moreover, the approximations for the subsets weighted with $\mu_{\infty}$ are not exact, since by Corollary 3.3 the sum over $\phi^{\text{SII}}(ij)$ is similar for $T=i$ and $T=N\setminus i$.
@@ 3.2 Optimal Higher-Order Approximations
Let $n\geq 2k$ and let the precision matrix $\mathbf{A}_{k}:=\lim_{\mu_{\infty}\to\infty}(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}$, then for $|S_{1}|,|S_{2}|=k$
$(\mathbf{A}_{k})_{S_{1}S_{2}}\overset{!}{=}\frac{(-1)^{k-|S_{1}\cap S_{2}|}}{n-k+1}\binom{n-k}{k-|S_{1}\cap S_{2}|}^{-1}.$
@@
Let $n\geq 2k$ and let $\mathcal{T}_{k}:=\{T\subseteq N:k\leq|T|\leq n-k\}$ be a set of subsets, which defines
@@
If $(\mathbf{Q}_{k})_{ST}:=(-1)^{s-|S\cap T|}m_{s}(t-|S\cap T|)$, where $m_{s}(t):=\frac{(n-s-t)!\cdot t!}{(n-s+1)!}$ from Definition 2.1, then
$\phi^{\text{SII}}_{k}\overset{!}{=}\mathbf{Q}_{k}\mathbf{y}^{-}_{k}+\lim_{\mu_{\infty}\to\infty}\operatorname*{arg\,min}_{\phi_{k}\in\mathbb{R}^{\binom{n}{k}}}\left\|\sqrt{\mathbf{W}_{k}}\left(\mathbf{y}^{+}_{k}-\mathbf{X}_{k}\phi_{k}\right)\right\|_{2}^{2}.$
Note that $\mathbf{Q}_{k}$ contains all SII weights (Fumagalli et al., 2023), which are required for the correct weighting of subsets in $T\notin\mathcal{T}_{k}$. This is necessary as, in contrast to $k=2$, the weights do not converge to the SII weights for $\mu_{\infty}\to\infty$ and $T\notin\mathcal{T}_{k}$, which requires to adjust the representation using $\mathbf{y}^{+}_{k}$ and $\mathbf{y}^{-}_{k}$. Presumably, different $\mu_{\infty}$ could resolve this issue, which we leave to future research.
@@ 3.3 KernelSHAP-IQ for Shapley Interactions
We introduce two variants of KernelSHAP-IQ for SII based on distinct optimization problems. Similar to KernelSHAP, we solve an approximated objective by sampling subsets from a distribution $p^{*}(T)\propto q(t)$ with sampling weights $q(t)\geq 0$ for $t=0,\dots,n$. We re-write Equation 1 as
$\phi^{*}=\lim_{\mu_{\infty}\to\infty}\operatorname*{arg\,min}_{\phi_{k}\in\mathbb{R}^{\binom{n}{k}}}\mathbb{E}_{T\sim p^{*}}\left[\left\|\sqrt{\mathbf{W}^{*}_{k}}\left(\mathbf{y}_{k}-\mathbf{X}_{k}\phi_{k}\right)\right\|_{2}^{2}\right],$
where we set $\mu_{\infty}\gg 1$ to account for the limit and adjust $(\mathbf{W}^{*}_{k})_{TT}:=(\mathbf{W}_{k})_{TT}/p^{*}(T)$. This expectation is then approximated using Monte Carlo sampling with $T\sim p^{*}$. Lastly, KernelSHAP-IQ estimates are computed by solving the approximated WLS objective with $\hat{\mathbf{W}}_{k},\hat{\mathbf{X}}_{k}$ and $\hat{\mathbf{y}}_{k}$, containing sampled subsets $T\subseteq N$ at each row. More details and pseudocode can be found in Appendix B.
Inconsistent KernelSHAP-IQ. Our baseline method, inconsistent KernelSHAP, extends on a method for the SV, known as $k_{\text{ADD}}$-SHAP (Pelegrina et al., 2023). Their approach utilizes the KernelSHAP weights and optimizes for all interactions up to order $k$ simultaneously via the weighting in Proposition 3.2, i.e. in total $\sum_{\ell=1}^{k}\binom{n}{\ell}$ interactions. Introducing the stacked matrices $\mathbf{X}_{\leq k}:=(\mathbf{X}_{1},\dots,\mathbf{X}_{k})$ and $\phi_{\leq k}:=[\phi_{1},\dots,\phi_{k}]^{T}$, this approach solves
@@
Including these higher-order terms improved the approximation of the SV and provably yields the exact SV for $k=n$ (Pelegrina et al., 2023). With inconsistent KernelSHAP, we propose to use the higher-order terms as estimates for SII.
@@
Interestingly, we observe empirically that solving Equation 3 yields the exact SVs as well for $k<n$.
@@
FSI is related to Equation 3 with different weights $\mathbf{X}_{\leq k}$, which optimize $k$-SII $\Phi_{k}$ instead of SII $\phi^{\text{SII}}_{\leq k}$.
@@
2:  $\mathbf{y}_{1}\leftarrow[\nu(T_{1}),\dots,\nu(T_{b})]^{T}$
@@
6:     end for
7:     $(\hat{\mathbf{W}}^{*}_{\ell})_{TT}\leftarrow\mu_{\ell}(t)\cdot w_{T}$triangleright weight adjustment
@@
11:        $(\mathbf{Q}_{\ell})_{ST}\leftarrow\mathbf{1}_{T\notin\mathcal{T}_{\ell}}\cdot\textsc{{SubsetWeightSII}}(T,S)$
12:        $(\hat{\mathbf{y}}^{+}_{\ell})_{T},(\hat{\mathbf{y}}^{-}_{\ell})_{T}\leftarrow\mathbf{1}_{T\in\mathcal{T}_{\ell}}(\hat{\mathbf{y}}_{\ell})_{T},\mathbf{1}_{T\notin\mathcal{T}_{\ell}}(\hat{\mathbf{y}}_{\ell})_{T}$
13:        $\hat{\phi}_{\ell}\leftarrow\mathbf{Q}_{\ell}\hat{\mathbf{y}}^{-}_{\ell}+\textsc{{SolveWLS}}(\hat{\mathbf{X}}_{\ell},\hat{\mathbf{y}}^{+}_{\ell},\hat{\mathbf{W}}^{*}_{\ell})$
14:     end if
15:     $\hat{\mathbf{y}}_{\ell+1}\leftarrow\hat{\mathbf{y}}_{\ell}-\hat{\mathbf{X}}_{\ell}\cdot\hat{\phi}_{\ell}$triangleright compute residuals
16:  end for
17:  $\hat{\Phi}_{k}\leftarrow\textsc{{AggregateSII}}(\hat{\phi}_{1},\dots,\hat{\phi}_{k})$triangleright compute $k$-SII
18:  return $k$-SII estimates $\hat{\Phi}_{k}$, SII estimates $\hat{\phi}_{1},\dots,\hat{\phi}_{k}$
KernelSHAP-IQ. We now introduce KernelSHAP-IQ based on the novel representation in Theorem 3.7, which ensures that KernelSHAP-IQ converges to SII for $k=2$. Further, with Conjecture 3.9, KernelSHAP-IQ can be extended to $k>2$, where we empirically validate its convergence. KernelSHAP-IQ iteratively constructs $\hat{\nu}_{\ell}$ for $\ell=1,\dots,k$. Starting from $\ell=1$, where KernelSHAP-IQ reduces to KernelSHAP, we estimate $\hat{\phi}^{\text{SII}}_{\ell}$ and compute the residuals $\hat{\mathbf{y}}_{\ell+1}:=\hat{\mathbf{y}}_{\ell}-\mathbf{X}_{\ell}\cdot\hat{\phi}^{\text{SII}}_{\ell}$ for the next iteration. At each step, we adjust $\mu_{\ell}$ by $w_{T}$, which accounts for the sampling probabilities $p^{*}(T)$ and the number of Monte Carlo samples. As a default, we propose the KernelSHAP weights for sampling, i.e $p(T)\propto\mu_{1}(t)$, and apply the border-trick (Fumagalli et al., 2023). For further details, we refer to Section B.3. We repeat this process until $\ell=k$, summarized in Algorithm 1. Lastly, the SII estimates are aggregated via the non-recursive formula (cf. Section B.6) to $k$-SII values, and used for final interpretation.

### v2

@@ 3.1 Optimal Approximations via Shapley Interactions
We first link $\hat{\nu}_{k}$ and SII, which reveals a surprising recursion.
@@
If $k\geq 2$, then
$\hat{\nu}_{k}(T)=\hat{\nu}_{k-1}(T)+\sum_{S\subseteq N}^{|S|=k}\phi^{\text{SII}}(S)\lambda(|S|,|S\cap T|)$
@@
It holds $\lambda(k,k)=0$ for $k>1$, $\hat{\nu}_{1}(T)=\sum_{i\in T}\phi^{\text{SV}}(i)$, and $\hat{\nu}_{2}(T)=\hat{\nu}_{1}(T)-\frac{1}{2}\sum_{{ij}\subseteq N}^{|ij\cap T|=1}\phi^{\text{SII}}(ij)$.
Increasing $k$ successively builds on the SV-based approximation $\hat{\nu}_{1}(T)$ from Equation 1 by including all SII of order $k$, i.e. $\binom{n}{k}$ components. With increasing $k$, the approximation $\hat{\nu}_{k}$ includes more interaction terms, which yields a complexity-accuracy trade-off. In this context, $\hat{\nu}_{1}$ based on the SVs is the least complex, and $\hat{\nu}_{n}$ is the most complex and most faithful approximation, where $\hat{\nu}_{n}(T)=\nu(T)$ follows from a game-theoretic result (Grabisch et al., 2000).
@@
For any $T\subseteq N$, it holds $\nu(T)=\hat{\nu}_{n}(T)$.
In this case, $n$-SII values ($k=n$ for $k$-SII), which construct $\hat{\nu}_{n}$, are the Möbius transform (Bordt & von Luxburg, 2023), which is related to the well-known functional ANOVA decomposition (Hooker, 2004) from statistics, if $\nu$ are conditional expectations (Herren & Hahn, 2022).
Optimality of the SII.
We now link SII of order $k$, $\phi^{\text{SII}}_{k}$, to an optimal $k$-additive approximation of $\nu$ in terms of a WLS objective. More concretely, we show that for all interaction indices $\phi_{k}$ of order $k$ with induced approximation via Proposition 3.2, the SII $\phi_{k}^{\text{SII}}$ yields the optimal approximation given a specific WLS objective. In the following, it is convenient to use matrix notations, where the rows correspond to subsets $T\subseteq N$ and the columns to interactions $S\subseteq N$ of order $|S|=k$. We first introduce the residual vector $\mathbf{y}_{k}\in\mathbb{R}^{2^{n}}$ as $(\mathbf{y}_{1})_{T}:=\nu(T)$ and for $k\geq 2$
$(\mathbf{y}_{k})_{T}:=\nu(T)-\hat{\nu}_{k-1}(T)\text{ for }T\subseteq N.$
Next, we introduce the coefficient matrix $\mathbf{X}_{k}\in\mathbb{R}^{2^{n}\times\binom{n}{k}}$ as
$(\mathbf{X}_{k})_{TS}:=\lambda(|S|,|T\cap S|)\text{ for }T,S\subseteq N\text{ with }|S|=k.$
With SII of order $k$, $\phi^{\text{SII}}_{k}\in\mathbb{R}^{\binom{n}{k}}$, and Proposition 3.2 the approximation error of $\nu$ and $\hat{\nu}_{k}$ is given by $\mathbf{y}_{k}-\mathbf{X}_{k}\phi^{\text{SII}}_{k}$. Introducing a diagonal weight matrix $\mathbf{W}_{k}\in\mathbb{R}^{2^{n}\times 2^{n}}$, we aim to characterize $\phi^{\text{SII}}_{k}$ as the solution to a WLS problem
@@
where $\|\cdot\|_{2}$ is the Euclidean norm. Note that this WLS problem for $k=1$ is similar to Equation 1 but includes all subsets $T\subseteq N$. The solution is explicitly given as
@@
We now define the weights of $\mathbf{W}_{k}$ with $\mu_{\infty}\gg 1$ as
@@
The weights $\mu_{k}$ appear naturally from $\phi^{\text{SII}}(S)=\sum_{T\subseteq N}\nu(T)\omega_{s}(t,|T\cap S|)$ (Fumagalli et al., 2023) as the common factors $\mu_{k}(t)\propto\omega_{s}(t,0),\dots\omega_{s}(t,k)$ for $k\leq t\leq n-k$. Since $(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1}\in\mathbb{R}^{\binom{n}{k}\times\binom{n}{k}}$ does not depend on $T$, it holds that $\mu_{k}$ controls the final weight in $\mathbf{X}_{k}^{T}\mathbf{W}_{k}$ for all interactions $S$. For a more detailed discussion, see Section B.9.
For $k=1$, $\mu_{1}$ are the KernelSHAP weights and the SV is the solution of the WLS problem for $\mu_{\infty}\to\infty$, which justifies KernelSHAP.
@@
A large weight $\mu_{\infty}$ requires the solution to achieve lower approximation error for this subset, which is a soft constraint. The subsets with weights $\mu_{\infty}$ are the empty set and $N$. For the empty set, the approximation error is constant, and is thus not influenced by the solution of the WLS problem. For $N$, the approximation error is $(\mathbf{y}_{1})_{N}-(\mathbf{X}_{1}\phi)_{N}=\nu(N)-\sum_{i\in N}\phi(i)$, which is zero if efficiency holds. Intuitively, with $\mu_{\infty}\to\infty$, the soft constraint becomes a hard constraint, which requires zero approximation error, i.e. efficiency with $\nu(N)=\sum_{i\in N}\phi(i)$. This was argued in KernelSHAP (Lundberg & Lee, 2017) without formal proof. Theorem 3.6 is a formal proof that this intuition holds, and indeed yields the SV satisfying the efficiency constraint from Equation 1.
For pairwise SII, we now present a novel representation akin to Theorem 3.6.
@@
While $\mu_{\infty}\to\infty$ corresponds to a constraint for $k=1$ that ensures efficiency, its behavior for $k=2$ is less clear. In fact, $\mu_{2}(1)=\mu_{2}(n-1)=\mu_{\infty}$, but for two such subsets $T\in\{\ell,N\setminus\ell\}$ with $\ell\in N$ the approximation is for both the same value as $(X_{2}\phi_{2})_{T}=-\frac{1}{2}\sum_{ij\subseteq N}^{|T\cap ij|=1}\phi(ij)=-\frac{1}{2}\sum_{j\in N}^{j\neq\ell}\phi(j\ell)$, which implies that the corresponding hard constraint of zero approximation error cannot be satisfied. Consequently, contrary to $k=1$, the soft constraints with $\mu_{\infty}$ cannot be re-written to a hard-constrained WLS problem, since it would not have a solution in general.
@@ 3.2 Optimal Higher-Order Approximations
Let $n\geq 2k$ and define the precision matrix
$\mathbf{A}_{k}:=\lim_{\mu_{\infty}\to\infty}(\mathbf{X}^{T}_{k}\mathbf{W}_{k}\mathbf{X}_{k})^{-1},$
Then, for $S_{1},S_{2}\subseteq N$ with $|S_{1}|=|S_{2}|=k$, it holds
$(\mathbf{A}_{k})_{S_{1}S_{2}}=\frac{(-1)^{k-|S_{1}\cap S_{2}|}}{n-k+1}\binom{n-k}{k-|S_{1}\cap S_{2}|}^{-1}.$
@@
Let $n\geq 2k$ and $\mathcal{T}_{k}:=\{T\subseteq N:k\leq|T|\leq n-k\}$, which splits for $T\subseteq N$ the vector $\mathbf{y}_{k}$ into
@@
Let further for $T\subseteq N$ and interaction $S\subseteq N$ with $|S|=k$
$(\mathbf{Q}_{k})_{ST}:=\begin{cases}(-1)^{s-|S\cap T|}m_{s}(t-|S\cap T|),&\text{if }T\notin\mathcal{T}_{k},\\ 0,&\text{if }T\in\mathcal{T}_{k},\end{cases}$
where $m_{s}(t):=\frac{(n-s-t)!\cdot t!}{(n-s+1)!}$ from Definition 2.1. Then,
$\phi^{\text{SII}}_{k}=\mathbf{Q}_{k}\mathbf{y}^{-}_{k}+\lim_{\mu_{\infty}\to\infty}\operatorname*{arg\,min}_{\phi_{k}\in\mathbb{R}^{\binom{n}{k}}}\left\|\sqrt{\mathbf{W}_{k}}\left(\mathbf{y}^{+}_{k}-\mathbf{X}_{k}\phi_{k}\right)\right\|_{2}^{2}.$
3.9 states that higher-order SII are again represented as the solution to a WLS problem. However, in contrast to $k\leq 2$, the representation does only hold partially for $T\in\mathcal{T}_{k}$. For $T\notin\mathcal{T}_{k}$, $\mathbf{Q}_{k}$ contains all SII weights (Fumagalli et al., 2023), which are required for the correct weighting. This is necessary as, in contrast to $k\leq 2$, the weights do not converge to the SII weights for $\mu_{\infty}\to\infty$ and $T\notin\mathcal{T}_{k}$, which requires to adjust the representation using $\mathbf{y}^{+}_{k}$ and $\mathbf{y}^{-}_{k}$. Presumably, different $\mu_{\infty}$ could resolve this issue, which we leave to future research.
@@ 3.3 KernelSHAP-IQ for Shapley Interactions
We introduce two variants of KernelSHAP-IQ for SII based on distinct optimization problems. Inconsistent KernelSHAP-IQ solves a single WLS problem and does not converge to SII. In contrast, (consistent) KernelSHAP-IQ solves iteratively a WLS problem for every order and converges to SII. Similar to KernelSHAP, we solve an approximated objective by sampling subsets from a distribution $p^{*}(T)\propto q(t)$ with sampling weights $q(t)\geq 0$ for $t=0,\dots,n$. We re-write Equation 2 as
$\phi^{*}=\operatorname*{arg\,min}_{\phi_{k}\in\mathbb{R}^{\binom{n}{k}}}\mathbb{E}_{T\sim p^{*}}\left[\left\|\sqrt{\mathbf{W}^{*}_{k}}\left(\mathbf{y}_{k}-\mathbf{X}_{k}\phi_{k}\right)\right\|_{2}^{2}\right],$
where we adjusted $(\mathbf{W}^{*}_{k})_{TT}:=(\mathbf{W}_{k})_{TT}/p^{*}(T)$. According to our previous results, the right-hand side has to be solved for $\mu_{\infty}\to\infty$, and we thus set $\mu_{\infty}\gg 1$ to account for the limit. KernelSHAP-IQ estimates are computed by solving the approximated WLS objective with $\hat{\mathbf{W}}_{k},\hat{\mathbf{X}}_{k}$ and $\hat{\mathbf{y}}_{k}$, containing sampled subsets $T\subseteq N$ at each row. More details and pseudocode can be found in Appendix B.
Inconsistent KernelSHAP-IQ. Our baseline method, inconsistent KernelSHAP, solves a single WLS problem using the KernelSHAP weights. We include all interactions up to order $k$ via the weighting in Proposition 3.2, i.e. in total $\sum_{\ell=1}^{k}\binom{n}{\ell}$ interactions. More formally, introducing the stacked matrices $\mathbf{X}_{\leq k}:=(\mathbf{X}_{1},\dots,\mathbf{X}_{k})$ and $\phi_{\leq k}:=[\phi_{1},\dots,\phi_{k}]^{T}$, this approach solves
@@
Inconsistent KernelSHAP solves the WLS problem with the approximated objective to obtain estimates for SII.
@@
Interestingly, we observe empirically that solving Equation 4 yields the exact SVs.
@@
Inconsistent KernelSHAP-IQ is related, but not equivalent, to $k_{\text{ADD}}$-SHAP (Pelegrina et al., 2023) and FSI. $k_{\text{ADD}}$-SHAP includes $\phi^{\text{SII}}(\emptyset)$ and a sum starting from zero in $\lambda$ in $\mathbf{X}_{\leq k}$. FSI uses binary weights $\mathbf{X}_{\leq k}$, and effectively solves for $k$-SII $\Phi_{k}$ instead of SII $\phi^{\text{SII}}_{\leq k}$.
@@
2:  $\hat{\mathbf{y}}_{1}\leftarrow[\nu(T_{1}),\dots,\nu(T_{b})]^{T}$
@@
6:        $(\hat{\mathbf{W}}^{*}_{\ell})_{TT}\leftarrow\mu_{\ell}(t)\cdot w_{T}$triangleright weight adjustment
7:     end for
@@
11:        for $T\in\{T_{i}\}$ and $|S|=\ell$ do
12:           $(\mathbf{Q}_{\ell})_{ST}\leftarrow\mathbf{1}_{T\notin\mathcal{T}_{\ell}}\cdot\textsc{{SIIWeight}}(T,S)$
13:           $(\hat{\mathbf{y}}^{+}_{\ell})_{T}\leftarrow\mathbf{1}_{T\in\mathcal{T}_{\ell}}(\hat{\mathbf{y}}_{\ell})_{T}$
14:           $(\hat{\mathbf{y}}^{-}_{\ell})_{T}\leftarrow\mathbf{1}_{T\notin\mathcal{T}_{\ell}}(\hat{\mathbf{y}}_{\ell})_{T}$
15:        end for
16:        $\hat{\phi}_{\ell}\leftarrow\mathbf{Q}_{\ell}\hat{\mathbf{y}}^{-}_{\ell}+\textsc{{SolveWLS}}(\hat{\mathbf{X}}_{\ell},\hat{\mathbf{y}}^{+}_{\ell},\hat{\mathbf{W}}^{*}_{\ell})$
17:     end if
18:     $\hat{\mathbf{y}}_{\ell+1}\leftarrow\hat{\mathbf{y}}_{\ell}-\hat{\mathbf{X}}_{\ell}\cdot\hat{\phi}_{\ell}$triangleright compute residuals
19:  end for
20:  $\hat{\Phi}_{k}\leftarrow\textsc{{AggregateSII}}(\hat{\phi}_{1},\dots,\hat{\phi}_{k})$triangleright compute $k$-SII
21:  return $k$-SII estimates $\hat{\Phi}_{k}$, SII estimates $\hat{\phi}_{1},\dots,\hat{\phi}_{k}$
KernelSHAP-IQ. We now introduce KernelSHAP-IQ based on the novel representation in Theorem 3.7, which ensures that KernelSHAP-IQ converges to SII for $k=2$. Further, with Conjecture 3.9, KernelSHAP-IQ can be extended to $k>2$, where we empirically validate its convergence. KernelSHAP-IQ once samples subsets given the sampling weights $q$, where the weights $w$ account for the sampling probabilities and the number of Monte Carlo samples (line 1). As a default, we propose the KernelSHAP weights for sampling, i.e $p(T)\propto\mu_{1}(t)$, and apply the border-trick (Fumagalli et al., 2023). For further details regarding the sampling and the weights $w$, we refer to Section B.3. The game is then evaluated on all sampled subsets, stored in $\hat{\mathbf{y}}_{1}$, which determines the computational complexity (line 2). Then, $\hat{\nu}_{\ell}$ is iteratively constructed for $\ell=1,\dots,k$ by computing the SII estimates (line 3-19). Starting from $\ell=1$, where KernelSHAP-IQ reduces to KernelSHAP, $\hat{\phi}_{\ell}$ is estimated and the residual $\hat{\mathbf{y}}_{\ell+1}:=\hat{\mathbf{y}}_{\ell}-\mathbf{X}_{\ell}\cdot\hat{\phi}_{\ell}$ is computed for the next iteration (line 18). We repeat this process until $\ell=k$, summarized in Algorithm 1. At each step, we set the matrices for the WLS problem (line 4-7). The coefficient matrix $\hat{\mathbf{X}}_{\ell}$ is set for every sampled subset and every interaction of the current order $\ell$ (line 5). The diagonal weight matrix $\hat{\mathbf{W}}^{*}_{\ell}$ is set for every sampled subset, and contains $\mu_{\ell}$ adjusted by $w_{T}$, which accounts for the sampling probabilities $p^{*}(T)$ and the number of Monte Carlo samples (line 6). If $\ell\leq 2$, then the WLS problem is solved directly with $\hat{\mathbf{X}}_{\ell},\hat{\mathbf{y}}_{\ell}$, and $\hat{\mathbf{W}}^{*}_{\ell}$ (line 9). If $\ell>2$, then Conjecture 3.9 applies, and the WLS problem is split by $\mathcal{T}_{\ell}$ (line 11-15). For every sampled subset and every interaction of the current order $\ell$, the SII weights are assigned to $\mathbf{Q}_{\ell}$ (line 12), and the residuals $\hat{\mathbf{y}}_{\ell}$ are split by $\mathcal{T}_{k}$ into $\hat{\mathbf{y}}^{+}_{\ell}$ (line 13) and $\hat{\mathbf{y}}^{-}_{\ell}$ (line 14). The SII estimates $\hat{\phi}_{\ell}$ are then computed with the SII weights $\mathbf{Q}_{\ell}$ for $\hat{\mathbf{y}}_{\ell}^{-}$ and by solving the WLS problem with $\hat{\mathbf{X}}_{\ell},\hat{\mathbf{y}}^{+}_{\ell}$ and $\hat{\mathbf{W}}^{*}_{\ell}$ (line 16). After reaching $\ell=k$, the SII estimates are aggregated via the non-recursive formula (cf. Section B.6) to $k$-SII values, and used for final interpretation (line 20-21).